# SeqAfford: Sequential 3D Affordance Reasoning via Multimodal Large Language Model

### v2

@@ Abstract
3D affordance segmentation aims to link human instructions to touchable regions of 3D objects for embodied manipulations. Existing efforts typically adhere to single-object, single-affordance paradigms, where each affordance type or explicit instruction strictly corresponds to a specific affordance region and are unable to handle long-horizon tasks. Such a paradigm cannot actively reason about complex user intentions that often imply sequential affordances. In this paper, we introduce the Sequential 3D Affordance Reasoning task, which extends the traditional paradigm by reasoning from cumbersome user intentions and then decomposing them into a series of segmentation maps. Toward this, we construct the first instruction-based affordance segmentation benchmark that includes reasoning over both single and sequential affordances, comprising 180K instruction-point cloud pairs. Based on the benchmark, we propose our model, SeqAfford, to unlock the 3D multi-modal large language model with additional affordance segmentation abilities, which ensures reasoning with world knowledge and fine-grained affordance grounding in a cohesive framework. We further introduce a multi-granular language-point integration module to endow 3D dense prediction. Extensive experimental evaluations show that our model excels over well-established methods and exhibits open-world generalization with sequential reasoning abilities. Code and Dateset are available at: https://github.com/hq-King/SeqAfford.
@@ 1 Introduction
Affordance is a crucial lens through which humans and embodied agents interact with various objects of the world. When provided with human instructions, affordance aims to highlight the actionable possibilities of these objects, linking visual perception with manipulation. While 2D affordance offers visual cues that suggest potential actions to embodied systems, 3D affordance provides a more direct and intuitive guidance for executing tasks in the realistic 3D world, and thus solidify the foundation for downstream robot manipulation tasks.
@@
Unfortunately, current systems are incapable of actively reasoning based on complex user intentions, breaking them down into actionable primitives, and formulating a chain of affordances derived from each primitive. Real-world physical interactions, in particular, require modeling the inherent complexity of human instructions that often involve handling multiple objects and navigating through sequential affordances. For example, when agents are required to “use the microwave to reheat the food inside the bowl ”, they must first grasp the bowl, then open the microwave before using it to contain the bowl. From this perspective, the multi-object sequential reasoning ability is indispensable for shaping the next-generation affordance systems.
@@
    We are the first to leverage the internalized world knowledge of the pre-aligned 3D MLLMs to achieve multi-target sequential reasoning and explanations in a cohesive framework.
@@ 2 Related Work
3D Affordance Segmentation. With the rapid advancement of embodied AI, research on 3D affordance has increasingly garnered significant attention from both academia and industry. 3D AffordanceNet [5], built on point cloud data from PartNet [21], was the first to construct a fine-grained 3D affordance dataset, establishing a benchmark for 3D affordance research. Building on this,  Yang et al. [39] proposed leveraging universal knowledge extracted from 2D human-object interaction (HOI) images to assist in 3D affordance segmentation. However, these methods rely solely on visual information to infer affordances, without considering that embodied agents in the real world communicate with humans through language. Consequently, such methods are limited in their applicability for direct deployment in embodied agents. Recently,  Li et al. [12] introduced a language-based affordance segmentation task, promoting the integration of natural language and affordance understanding.
@@ 4.2 Network Architecture
To enable multi-granular feature propagation, as shown in Fig. 4, we hierarchically up-sample (UpS. in Fig. 4) the intermediate features from the 3D encoder, and propagate features through farthest point sampling (FPS in Fig. 4) process to sequentially generate ${f}_{1}$, ${f}_{2}$, and the ultimate dense feature ${f}_{\text{dense}}$. During the up-sampling process of intermediate features, various feature up-sample techniques are adopted, inspired by PointNet++ [26] and DGCNN [34]. More details about the multi-granular feature propagation process are revealed in the supplementary details.
During the point-language integration stage, the model takes the dense point cloud features ${f}_{\text{dense}}$, the sparse point cloud features ${f}_{\text{sparse}}$ and the instruction-rich $\bf H^{(i)}_{\text{seg}}$ as input. The ${\bf H}^{(i)}_{\text{seg}}$ and dense features ${f}_{\text{dense}}$ are used as $Q$ and $K$, $V$ respectively to perform cross-attention, and FFN, and the results obtained are fused with the sparse features ${f}_{\text{sparse}}$ to get $\bf A_{f}$. Finally, the decoder takes $A_{f}$ as the input to get affordance mask $\tilde{\bf y}_{\text{mask}}$
$A_{f}$ $=\mathcal{G}({f}_{\text{dense}},{f}_{\text{sparse}},H_{\text{seg}}),\quad\tilde{y}_{\text{mask}}=\mathcal{D}({\bf A}_{f}),$ (3)

### v3

@@ Abstract
3D affordance segmentation aims to link human instructions to touchable regions of 3D objects for embodied manipulations. Existing efforts typically adhere to single-object, single-affordance paradigms, where each affordance type or explicit instruction strictly corresponds to a specific affordance region and are unable to handle long-horizon tasks. Such a paradigm cannot actively reason about complex user intentions that often imply sequential affordances. In this paper, we introduce the Sequential 3D Affordance Reasoning task, which extends the traditional paradigm by reasoning from cumbersome user intentions and then decomposing them into a series of segmentation maps. Toward this, we construct the first instruction-based affordance segmentation benchmark that includes reasoning over both single and sequential affordances, comprising 180K instruction-point cloud pairs. Based on the benchmark, we propose our model, SeqAfford, to unlock the 3D multi-modal large language model with additional affordance segmentation abilities, which ensures reasoning with world knowledge and fine-grained affordance grounding in a cohesive framework. We further introduce a multi-granular language-point integration module to endow 3D dense prediction. Extensive experimental evaluations show that our model excels over well-established methods and exhibits open-world generalization with sequential reasoning abilities. Project page: https://seq-afford.github.io/.
@@ 1 Introduction
Affordance is a crucial lens through which humans and embodied agents interact with various objects of the world . When provided with human instructions, affordance aims to highlight the actionable possibilities of these objects, linking visual perception with manipulation. While 2D affordance [35, 14, 47] offers visual cues that suggest potential actions to embodied systems, 3D affordance [6, 3, 51, 8, 24] provides a more direct and intuitive guidance for executing tasks in the realistic 3D world, and thus solidify the foundation for downstream robot manipulation tasks [45, 30, 11, 44].
@@
Current systems can’t actively reason complex user intentions, break them down into actionable primitives, or formulate chains of affordances from each primitive. Real-world physical interactions require modeling complex human instructions involving multiple objects and sequential affordances. For example, to ’reheat food in a bowl using the microwave’, agents must first grasp the bowl, open the microwave, then contain the bowl inside. Thus, multi-object sequential reasoning is crucial for next-generation affordance systems.
@@
    We leverage the internalized world knowledge of pre-aligned 3D MLLMs to achieve fine-grained affordance-level multi-target sequential reasoning and explanations in a cohesive framework.
@@ 2 Related Work
3D Affordance Segmentation. With the rapid advancement of embodied AI, research on 3D affordance has increasingly garnered significant attention from both academia and industry. 3D AffordanceNet [7], built on point cloud data from PartNet [26], was the first to construct a fine-grained 3D affordance dataset, establishing a benchmark for 3D affordance research. Building on this, some researchers  [50, 38] proposed leveraging universal knowledge extracted from 2D human-object interaction (HOI) images to assist in 3D affordance segmentation. However, these methods rely solely on visual information to infer affordances, without considering that embodied agents in the real world communicate with humans through language. Consequently, such methods are limited in their applicability for direct deployment in embodied agents. Recently,  Li et al. [16] introduced a language-based affordance segmentation task, promoting the integration of natural language and affordance understanding.
@@ 4.2 Network Architecture
To enable multi-granular feature propagation, as shown in Fig. 4, we hierarchically up-sample (UpS. in Fig. 4) the intermediate features from the 3D encoder, and propagate features through farthest point sampling (FPS in Fig. 4) process to sequentially generate ${\bf f}_{1}$, ${\bf f}_{2}$, and the ultimate dense feature ${\bf f}_{\text{dense}}$. During the up-sampling process of intermediate features, various feature up-sample techniques are adopted, inspired by PointNet++ [32] and DGCNN [42]. More details about the multi-granular feature propagation process are revealed in the supplementary details.
During the point-language integration stage, the model takes the dense point cloud features ${\bf f}_{\text{dense}}$, the sparse point cloud features ${\bf f}_{\text{sparse}}$ and the instruction-rich $\bf H^{(i)}_{\text{seg}}$ as input. The ${\bf H}^{(i)}_{\text{seg}}$ and dense features ${\bf f}_{\text{dense}}$ are used as $\bf Q$ and $\bf K$, $\bf V$ respectively to perform cross-attention, and FFN, and the results obtained are fused with the sparse features ${\bf f}_{\text{sparse}}$ to get ${\bf A}^{(i)}_{f}$. Finally, the decoder takes ${\bf A}^{(i)}_{f}$ as the input to get affordance mask $\tilde{\bf y}_{\text{mask}}$
${\bf A}^{(i)}_{f}$ $=\mathcal{G}({\bf f}_{\text{dense}},{\bf f}_{\text{sparse}},{\bf H}^{(i)}_{\text{seg}}),\quad\tilde{y}_{\text{mask}}=\mathcal{D}({\bf A}^{(i)}_{f}),$ (3)
@@ 6 Conclusion
7 Acknowledgement
This work was supported by the Shanghai Local College Capacity Building Program (23010503100), National Natural Science Foundation of China (No.62406195), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), HPC Platform of ShanghaiTech University and Shanghai Engineering Research Center of Intelligent Vision and Imaging.